OpenStudy (anonymous):

Need Help! Find derivative of f(x)=1-1/2sqrt(1-x)

OpenStudy (anonymous):

I'm really confused on how to find the derivative of fraction square root part..do I do the quotient rule?

OpenStudy (anonymous):

oh no !!

OpenStudy (anonymous):

as much as i lectured before about not using rational exponents, that was when you had \[\frac{d}{dx}\sqrt{f(x)}\]

OpenStudy (astrophysics):

\[f(x) = 1-\frac{ 1 }{ 2\sqrt{1-x} }?\]

OpenStudy (anonymous):

Yep, that's it astrophysics!

OpenStudy (anonymous):

for this one you probably do want to use rational exponents

OpenStudy (anonymous):

LOL THANKS SATTELITE!

OpenStudy (anonymous):

\[1-\frac{1}{2}(1-x)^{-\frac{1}{2}}\]

OpenStudy (astrophysics):

ye was gonna do that

OpenStudy (anonymous):

then go nuts

OpenStudy (anonymous):

i guess you are looking for the second derivative of the previous one right?

OpenStudy (anonymous):

wow...so like why couldnt I use the rational exponents for the other question we were doing? and yea

OpenStudy (anonymous):

of course you could, what i meant was it is not necessary once you know it because after you use the power rule, you are going to have to convert out of the rational exponents to do any sort of computation

OpenStudy (anonymous):

i wasn't saying you couldn't use it, i was saying you SHOULDN'T use it because you should just memorize the derivative of the square root function or its composition

OpenStudy (anonymous):

kk gotcha

OpenStudy (anonymous):

this one would be a bear to do without the power rule, but let me lecture once again

OpenStudy (anonymous):



OpenStudy (astrophysics):

Power rule and chain rule

OpenStudy (anonymous):

DO NOT USE THE QUOTIENT RULE IF THE NUMERATOR IS A NUMBER !!

OpenStudy (astrophysics):

^ this, I see so many people applying quotient rule to problems like that

OpenStudy (anonymous):

i will pontificate more on that in a moment, lets finish this one

OpenStudy (anonymous):

Okay enough about bashing on mwah, let's finish the problem

OpenStudy (anonymous):

\[-\frac{1}{4}(1-x)^{-\frac{3}{2}}\] is your derivative right?

OpenStudy (anonymous):

Yea and then I solved it down to what I have above

OpenStudy (anonymous):

are you trying to set it equal to zero?

OpenStudy (anonymous):

yea...should I not

OpenStudy (anonymous):

well not you should not again the curse of rational exponents

OpenStudy (anonymous):

\[-\frac{1}{4\sqrt{(1-x)^2}}\]

OpenStudy (anonymous):

crap

OpenStudy (anonymous):

\[-\frac{1}{4\sqrt{(1-x)^3}}\]

OpenStudy (anonymous):

a fraction is only zero if the numerator is

OpenStudy (anonymous):

so forget about setting this equal to zero, it is never zero lesson: once you have the derivative (whatever method) get it out of exponential notation and write what it really is

OpenStudy (anonymous):

my guess is you are asked about concavity right?